# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v2

@@ Abstract.
High-Level Synthesis enables the rapid prototyping of hardware accelerators, by combining a high-level description of the functional behavior of a kernel with a set of micro-architecture optimizations as inputs. Such optimizations can be described by inserting pragmas for e.g. pipelining and replication of units, or even higher level transformations for HLS such as automatic data caching using the AMD/Xilinx Merlin compiler. Selecting the best combination of pragmas, even within a restricted set, remains particularly challenging and the typical state-of-practice uses design-space exploration to navigate this space. But due to the highly irregular performance distribution of pragma configurations, typical DSE approaches are either extremely time consuming, or operating on a severely restricted search space.
In this work we propose a framework to automatically insert HLS pragmas in regular loop-based programs, supporting pipelining, unit replication (coarse- and fine-grain), and data caching. We develop an analytical performance and resource model as a function of the input program properties and pragmas inserted, using non-linear constraints and objectives. We prove this model provides a lower bound on the actual performance after HLS. We then encode this model as a Non-Linear Program, by making the pragma configuration unknowns of the system, which is computed optimally by solving this NLP. This approach can also be used during DSE, to quickly prune points with a (possibly partial) pragma configuration, driven by lower bounds on achievable latency. We extensively evaluate our end-to-end, fully implemented system, showing it can effectively manipulate spaces of billions of designs in seconds to minutes for the kernels evaluated.
HLS, FPGA, Non-Linear Programming, Program Optimization
copyright: noneconference: ; ;
@@ 1. Introduction
High-level synthesis (HLS) (Cong et al., 2011; Zhang et al., 2008) compilers (Xilinx, 2023b; Intel, 2023; Siemens, 2023; Microchip, 2023) and source-to-source compiler for HLS (Xilinx, 2023a; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020) can reduce development time while delivering a good performance for the designs. However, achieving a satisfactory Quality of Results (QoR) often requires design-space exploration (DSE). This is because the design space, including which pragmas to insert and where, can not only contain millions of points, but typically does not present characteristics suitable for fast analytical exploration, such as convexity and regularity. Although the existing DSE methods (Sohrabizadeh et al., 2021, 2022; Wu et al., 2023) can find designs with a good QoR, it comes at a high computation cost: for example, hundreds of designs may be concretely instantiated using HLS to compute its estimated QoR during exploration (Sohrabizadeh et al., 2021).
@@
Our main objective in this work is to provide a system to automatically insert a set of hardware pragmas for HLS, that delivers a good QoR and yet significantly reduces the search time needed to obtain the final design. To address this challenge, we propose NLP-DSE , a framework built on top of the AMD/Xilinx Merlin source-to-source compiler (Xilinx, 2023a). This framework automatically inserts pragmas for unrolling/parallelization, pipelining, tiling and data caching for affine programs (Lattner et al., 2021), prior to HLS. These Merlin pragmas can also be inserted using a DSE approach, such as AutoDSE (Sohrabizadeh et al., 2021) or HARP (Sohrabizadeh et al., 2023a), which we use as reference for our evaluations. However, in contrast to AutoDSE (Sohrabizadeh et al., 2021) and HARP (Sohrabizadeh et al., 2023a), we specifically restrict the class of programs we manipulate to affine programs that are regular loop-based computations. In turn, it enables us to develop a hybrid analytical approach to drive the search, combined with a lightweight DSE to reduce the number of designs actually explored. NLP-DSE preserves, and often even improves, the final QoR of designs produced.
@@
To this end, we create a novel Non-Linear Programming approach to automatically insert pragmas in an existing program. We develop an analytical model combining latency and resources, targeting regular loop-based kernels (PolyBench, [n. d.]), that is parameterized by the pragma configuration. We can then designate the pragma configuration as the unknowns of this model, solving it by NLP to obtain the set of pragmas that minimizes latency. An important design principle of our approach is to ensure that the latency computed is a lower bound on the achievable latency for a given pragma configuration. This enables efficient pruning during DSE: any design predicted to have a latency lower bound higher than the best latency obtained through exploration so far is necessarily slower and does not need exploration. To overcome the fact that optimizing compilers, and the overall HLS toolchain underneath, may not apply optimizations as expected (e.g., due to insufficient resources, or limitations of the compiler’s implementation), we develop a lightweight NLP-based DSE approach, exploring parts of the design space with different types and amounts of hardware parallelism and array partitioning factors. We make the following contributions:
@@
  • We implement NLP-DSE it as an end-to-end, fully automated system and use it to conduct extensive evaluation on 47 benchmarks including kernels from linear algebra, image processing, physics simulation, graph analytics, datamining, etc. (PolyBench, [n. d.]). Our results show the ability of our approach to find in most cases a better QoR than AutoDSE, in significantly less time. Furthermore, in most instances, our approach outperforms HARP in terms of QoR within a comparable timeframe.
The paper is organized as follows. Section 2 motivates our approach and solution proposed. Section 3 presents our analytical performance and resource model. Section 4 delves into proving it is a lower bound on the final QoR. In Section 5, we introduce a non-linear formulation based on this model to automatically find pragma configurations by NLP optimization. Section 6 presents our lightweight DSE approach. Finally, sections 7 and 10 are devoted to evaluating our method validating the effectiveness of our approach and presenting related work, before concluding.
@@ 2.1. Pragma-based Optimizations for HLS
This work targets the automatic optimization of FPGA designs using HLS (Cong et al., 2011; Zhang et al., 2008), in particular when using compilers to automatically generate HLS-friendly optimized programs such as the AMD/Xilinx Merlin Compiler (Cong et al., 2016b, a; Xilinx, 2023a). Falcon Computing Solutions developed this source-to-source automation tool for FPGAs, which was acquired by Xilinx in late 2020 and is now open source.
HLS has made FPGA usage more accessible, and many projects are looking to further democratize this field by automating optimizations (Sohrabizadeh et al., 2021; Lai et al., 2019; Huang et al., 2021; Ye et al., 2022; Ferretti et al., 2018). Merlin was developed to improve the performance, reduce the development time of HLS-based designs and simplify the search space. To achieve this goal, it automatically generates data transfers between off-chip and on-chip memory and inserts important Vitis pragmas, such as for array partitioning. Moreover, Merlin does specific code transformation for the targeting hardware in adequacy with the hardware directive describe by the user. Additionally, it enables parallel coarse-grained processing elements by encapsulating inner loops as functions.
Merlin includes hardware directive pragmas such as:
@@
  • ACCEL pipeline <II=y> for pipelining;
@@
  • ACCEL cache <array=a> which transfers all required elements of array $a$ from off-chip to on-chip to perform computations within the specified sub-region. If the user does not specify this pragma, it can be applied automatically by Merlin.
@@
We note our approach is not restricted to Merlin, nor a particular version of a toolchain: by adjusting the parameters of the performance model, such as operation latency, resource usage per operation(s), etc. one can easily target other toolchains than the one we evaluate here.
@@ 2.2. DSE for Pragma Insertion
Design-space exploration techniques typically trade-off coverage for speed (Sohrabizadeh et al., 2021; Zuo et al., 2015b). That is, it may impose restrictions on the input programs supported, on the pragmas/transformations considered, etc. in order to accelerate the search (Zuo et al., 2015b; Pouchet et al., 2013a). To overcome the difficulty of providing accurate performance models for arbitrary programs, HLS may be invoked to obtain a QoR estimate, without imposing any restriction on the input programs and transformations used. However HLS time for highly optimized designs combining various Merlin pragmas (e.g., parallelism and caching) can quickly reach several hours per design, making the search process particularly time-consuming. One may restrict the space of pragmas considered, and especially their parameter range, to reduce search time. General-purpose DSE approaches such as AutoDSE (Sohrabizadeh et al., 2021) are by design agnostic to the input program features and search space explored, preserving generality. But as we demonstrate in this paper, it also misses opportunities for search acceleration that can be provided by careful static analysis, leading to missed performance opportunities.
@@
We illustrate the performance merits and limitations of such general-purpose DSE (Sohrabizadeh et al., 2021) on three important loop-based linear algebra benchmarks. GEMM, the classical dense general matrix-multiply, and 2mm shown in Listing 1 which computes the product of three matrices $D=alpha*A*B*C+beta*D$. Both are key computations in e.g., inference of transformers (Devlin et al., 2018). Gramschmidt computes QR decomposition using the Gram-Schmidt process. In later Section 7 we evaluate these benchmarks using various problem sizes, ranging from kBs to MBs of footprints for the matrices to demonstrate the robustness of our approach to varying and large problem sizes. Below we use matrices of about 300kB each.
1 Loop0: for (i1 = 0; i1 < 180; i1++)
@@ 11 S3: D[i2][j2] += tmp[i2][k2] * C[k2][j2];
13}
Listing 1: 2mm code: $D=\alpha\times A\times B\times C+\beta\times D$
@@
Table 1 displays the performance (in GigaFlop/s, GF/s) of the original programs, from PolyBench/C (PolyBench, [n. d.]), when fed to Merlin as-is. The best design found by AutoDSE, given a time budget of 20 hours per benchmark and a timeout of 3 hours per HLS run, is also reported. AutoDSE uses a bottleneck-driven search approach, which targets the improvement of the code section with the lowest throughput (Sohrabizadeh et al., 2021). It unambiguously achieves particularly solid improvements over a naive design without pragmas. However, we show in Table 3 below that a carefully built DSE technique, exploiting the regularity of affine programs and leveraging non-linear programming, can provide order(s) of magnitude higher performance for these exact benchmarks, all while using significantly less search time.
@@ 2.3. Limitations of General-Purpose DSE
AutoDSE (Sohrabizadeh et al., 2021) utilizes the HLS compilers as a black box, in order to select the configurations that minimize the objective function. The tools are agnostic of the input program shape and if AutoDSE detects that Merlin did not apply the pragmas as expected it allows the DSE to prune the design after Merlin has generated the HLS-C code. These frameworks use incremental DSEs, i.e., having no information on the characteristics of the program, they explore the space by increasing the parallelism in order to respond to a problem, e.g., a bottleneck for Auto-DSE.
@@
Table 2. Investigation into the design space and exploration extent concerning synthesized designs, pruned designs, and designs reaching timeout by the AutoDSE framework for the kernels 2mm, Gemm, and Gramschmidt.
Over Parallelization
@@
Botlleneck analysis will make it possible to select which part of the code to optimize as a priority (Sohrabizadeh et al., 2021). However, this priority does not take into account parallelism (i.e., hardware resources) that shall be deployed for other parts of the code. This creates code with regions that are extremely/fully parallelized, and others without any parallelism.
For 2mm, the fastest design found by AutoDSE mainly optimizes one loop body. When AutoDSE tries to optimize the second loop body, it favors the unroll factors to the power of two for the innermost loop and goes directly to the outermost loop. This does not improve performance or create configurations which are pruned. The fact that it does not try the other unrolls factors for the innermost loop before optimizing the other pragmas creates a loss of performance. For Gemm and Gramschmidt, the DSE finds designs with a good QoR. However the DSE wastes much time exploring too large unroll factors, which generates ponderously long synthesis times without giving any result as the HLS timeout is reached. The time spent increasing the unroll factor for certain pragmas without result does not allow the unroll factor of other pragmas to be increased, which results in missed performance.
2.4. Overview of NLP-DSE

### v4

@@ Abstract.
High-Level Synthesis enables the rapid prototyping of hardware accelerators, by combining a high-level description of the functional behavior of a kernel with a set of micro-architecture optimizations as inputs. Such optimizations can be described by inserting pragmas e.g. pipelining and replication of units, or even higher level transformations for HLS such as automatic data caching using the AMD/Xilinx Merlin compiler. Selecting the best combination of pragmas, even within a restricted set, remains particularly challenging and the typical state-of-practice uses design-space exploration to navigate this space. But due to the highly irregular performance distribution of pragma configurations, typical DSE approaches are either extremely time consuming, or operating on a severely restricted search space.
This work proposes a framework to automatically insert HLS pragmas in regular loop-based programs, supporting pipelining, unit replication, and data caching. We develop an analytical performance and resource model as a function of the input program properties and pragmas inserted, using non-linear constraints and objectives. We prove this model provides a lower bound on the actual performance after HLS. We then encode this model as a Non-Linear Program, by making the pragma configuration unknowns of the system, which is computed optimally by solving this NLP. This approach can also be used during DSE, to quickly prune points with a (possibly partial) pragma configuration, driven by lower bounds on achievable latency. We extensively evaluate our end-to-end, fully implemented system, showing it can effectively manipulate spaces of billions of designs in seconds to minutes for the kernels evaluated.
High-Level Synthesis, Field-programmable gate array, Non-Linear Programming, Program Optimization, Pragma insertion
copyright: rightsretainedjournal: TODAESjournalyear: 2025journalvolume: 30journalnumber: 2article: 26publicationmonth: 02doi: 10.1145/3711847ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
@@ 1. Introduction
High-level synthesis (HLS) (Cong et al., 2011; Zhang et al., 2008) compilers (Xilinx, 2023b; Intel, 2023; Siemens, 2023; Microchip, 2023) and source-to-source compiler for HLS (Xilinx, 2023a; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020) can reduce development time while delivering a good performance for the designs. However, achieving a satisfactory Quality of Results (QoR) often requires design-space exploration (DSE). This is because the design space, including which pragmas to insert and where, can not only contain millions of points, but typically does not present characteristics suitable for fast analytical exploration, such as convexity and regularity. Although the existing DSE methods (Sohrabizadeh et al., 2022b, a; Wu et al., 2023) can find designs with a good QoR, it comes at a high computation cost: for example, hundreds of designs may be concretely instantiated using HLS to compute its estimated QoR during exploration (Sohrabizadeh et al., 2022b). Alternatively, models can be built to estimate performance and resource usage instead of applying HLS (Zhao et al., 2017; Sohrabizadeh et al., 2023a). These approaches typically rely on enumerating the search space and estimating the QoR of candidate designs without deploying HLS, filtering out candidates with lower performance. They can quickly estimate the QoR of thousands of designs in a search space, for example using purely analytical models (Zhao et al., 2017), or deploying machine learning to learn performance predictors from large database (Bai et al., 2024) of designs with HLS synthesis results (Sohrabizadeh et al., 2023a; Qin et al., 2024; Ding et al., 2024). However, these approaches still face challenges in predicting the performance of new designs unseen from the training set, despite the recent success in transfer learning (Qin et al., 2024).
@@
Our main objective in this work is to provide a system to automatically insert a set of hardware pragmas for HLS, that delivers a good QoR and yet significantly reduces the search time needed to obtain the final design. To address this challenge, we propose NLP-DSE, a framework built on top of the AMD/Xilinx Merlin source-to-source compiler (Xilinx, 2023a). This framework automatically inserts pragmas for unrolling/parallelization, pipelining, tiling and data caching for affine programs (Lattner et al., 2021), prior to HLS. These Merlin pragmas can also be inserted using a DSE approach, such as AutoDSE (Sohrabizadeh et al., 2022b) or HARP (Sohrabizadeh et al., 2023a), which we use as reference for our evaluations. However, in contrast to AutoDSE (Sohrabizadeh et al., 2022b) and HARP (Sohrabizadeh et al., 2023a), we specifically restrict the class of programs we optimize to affine programs that are regular loop-based computations. In turn, it enables us to develop a hybrid analytical approach to drive the search, combined with a lightweight DSE to reduce the number of designs actually explored. NLP-DSE preserves, and often even improves, the final QoR of designs produced while significantly reducing exploration time.
@@
To this end, we create a novel Non-Linear Programming approach to automatically insert pragmas in an existing affine program. We develop an analytical model combining latency and resources, targeting regular loop-based kernels (Pouchet and Yuki, 2023), that is parameterized by the pragma configuration. We can then designate the pragma configuration as the unknowns of this model, solving it by NLP to obtain the set of pragmas that minimizes latency. An important design principle of our approach is to ensure that the latency computed is a lower bound on the achievable latency for a given pragma configuration. This enables efficient pruning during DSE: any design predicted to have a latency lower bound higher than the best latency obtained through exploration so far is necessarily slower and does not need exploration. To overcome the fact that optimizing compilers, and the overall HLS toolchain underneath, may not apply optimizations as expected (e.g., due to insufficient resources, or limitations of the compiler’s implementation), we develop a lightweight NLP-based DSE approach, exploring parts of the design space with different types and amounts of hardware parallelism and array partitioning factors. We make the following contributions:
@@
  • We implement NLP-DSE, an end-to-end, fully automated system and use it to conduct extensive evaluation on 47 benchmarks including kernels from linear algebra, image processing, physics simulation, graph analytics, datamining, etc. (Pouchet and Yuki, 2023). Our results show the ability of our approach to find in most cases a better QoR than AutoDSE, in significantly less time. Furthermore, in most instances, our approach outperforms HARP in terms of QoR within a comparable timeframe.
The paper is organized as follows. Section 2 motivates our approach and solution proposed. Section 4 presents our analytical performance and resource model. In Section 5, we introduce a non-linear formulation based on this model to automatically find pragma configurations by NLP optimization. Section 6 delves into proving it is a lower bound on the final QoR. Section 7 presents our lightweight DSE approach. Finally, sections 8 and 11 are devoted to evaluating our method validating the effectiveness of our approach and presenting related work, before concluding.
@@ 2.1. Pragma-based Optimizations for HLS
This work targets the automatic optimization of FPGA designs using HLS (Cong et al., 2011; Zhang et al., 2008), in particular when using compilers to automatically generate HLS-friendly optimized programs such as the AMD/Xilinx Merlin Compiler (Cong et al., 2016b, a; Xilinx, 2023a). Falcon Computing Solutions developed this source-to-source automation tool for FPGAs, which was acquired by Xilinx in 2020 and is now open source.
HLS has made FPGA usage more accessible, and many projects are looking to further democratize this field by automating optimizations (Sohrabizadeh et al., 2022b; Lai et al., 2019; Huang et al., 2021; Ye et al., 2022; Ferretti et al., 2018).
Merlin was developed to improve the performance, reduce the development time of HLS-based designs and simplify the search space. To achieve this, Merlin automatically generates data transfers between off-chip and on-chip memory based on the specified available on-chip memory and allows overlapping of memory transfers for different arrays when their transfers occur consecutively in the code. It also automatically inserts essential Vitis pragmas, such as array partitioning, based on the unrolling factor (pragma ACCEL parallel). Furthermore, Merlin performs hardware-specific code transformations in line with user-defined directives. For example, it applies strip-mining to loops that are partially unrolled, fully unrolling the newly created inner loop with a trip count equal to the user-specified unroll factor. Additionally, Merlin enables parallel coarse-grained processing by encapsulating inner loops as separate functions.
The pragmas available for use with Merlin are:
@@
  • ACCEL pipeline flatten <II=y> for pipelining;
@@
  • ACCEL cache <array=a> which transfers all required elements of array $a$ from off-chip to on-chip to perform computations within the specified sub-region. If the user does not specify this pragma, it can be applied automatically by Merlin;
  • ACCEL pipeline which creates a double buffer and enables overlapping of computation and communication. This pragma frequently fails to apply as expected in various scenarios, so we opted to exclude it from the design space to preserve a more accurate model. Incorporating it into the model requires a minor adjustment to the objective function.
@@
We note our approach is not restricted to Merlin, nor a particular version of a toolchain: by adjusting the parameters of the performance model, such as operation latency, resource usage per operation(s), etc. one can easily target other toolchains than the one we evaluate here. In a recent study, we employed a similar NLP-based approach directly on HLS programs (Pouget et al., 2025).
@@ 2.2. DSE for Pragma Insertion
Design-space exploration techniques typically trade-off coverage for speed (Sohrabizadeh et al., 2022b; Zuo et al., 2015b). That is, it may impose restrictions on the input programs supported, on the pragmas/transformations considered, etc. in order to accelerate the search (Zuo et al., 2015b; Pouchet et al., 2013). To overcome the difficulty of providing accurate performance models for arbitrary programs, HLS may be invoked to obtain a QoR estimate, without imposing any restriction on the input programs and transformations used. However HLS time for highly optimized designs combining various Merlin pragmas (e.g., parallelism and caching) can quickly reach tens of minutes to several hours per design, making the search process particularly time-consuming. One may restrict the space of pragmas considered, and especially their parameter range, to reduce search time. General-purpose DSE approaches, such as AutoDSE (Sohrabizadeh et al., 2022b), are agnostic to the input program features and the search space explored, thereby preserving generality. But as we demonstrate in this paper, it also misses opportunities for search acceleration that can be provided by careful static analysis, leading to missed performance opportunities.
@@
We illustrate the performance merits and limitations of such general-purpose DSE (Sohrabizadeh et al., 2022b) on three important loop-based linear algebra benchmarks. GEMM, the classical dense general matrix-multiply, and 2mm shown in Listing 1 which computes the product of three matrices $D=alpha*A*B*C+beta*D$. Both are key computations in e.g., inference of transformers (Devlin et al., 2019). Gramschmidt computes QR decomposition using the Gram-Schmidt process. In later Section 8 we evaluate these benchmarks using various problem sizes, ranging from kBs to MBs of footprints for the matrices to demonstrate the robustness of our approach to varying and large problem sizes. Below we use matrices of about 300kB each.
1Loop0: for (i1 = 0; i1 < 180; i1++)
@@ 11 S3: D[i2][j2] += tmp[i2][k2] * C[k2][j2];
Listing 1: 2mm code: $D=alpha\times A\times B\times C+beta\times D$
@@
AutoDSE
Table 1 displays the performance (in GigaFlop/s, GF/s) as reported by Vitis HLS 2021.1 estimation, targeting the AMD/Xilinx Alveo U200 FPGA. We report the performance of the original programs from PolyBench/C (Pouchet and Yuki, 2023) using the medium dataset size for 2mm and Gemm and large dataset for Gramschmidt, when fed to Merlin as-is (column Merlin). The best design found by AutoDSE, given a time budget of 20 hours per benchmark and a timeout of 3 hours per HLS run, is also reported. AutoDSE uses a bottleneck-driven search approach, which targets the improvement of the code section with the lowest throughput (Sohrabizadeh et al., 2022b). It unambiguously achieves particularly solid improvements over a naive design without pragmas. However, we show in Table 4 below that a carefully built DSE technique, exploiting the regularity of affine programs and leveraging non-linear programming, can provide order(s) of magnitude higher performance for these exact benchmarks, all while using significantly less search time.
@@
HARP
Machine learning can be deployed to accelerate the search space exploration, substituting for the actual HLS of a particular pragma configuration to estimate its QoR. Table 2 displays the performance achieved by HARP (Hierarchical Augmentation for Representation with Pragma optimization) (Sohrabizadeh et al., 2023a) on the same setup as in Table 1. The HARP model was trained on a database of about 5000 designs, comprising different kernels and problem sizes (including 2mm and gemm, but excluding gramschmidt), as well as different pragma configurations for these, following exactly the methodology in (Sohrabizadeh et al., 2023a). We evaluate 75,000 designs using HARP, and run HLS on the top-10 predicted designs. A timeout of 3 hours for each HLS run is implemented. We explicitly do not employ re-training or fine-tuning of the model for the 3 benchmarks reported, which requires additional HLS runs, and study the robustness of the model to varying affine benchmarks and problem sizes. We show in Table 4 below our proposed NLP-DSE approach can match and often outperform HARP, using less HLS runs that would typically be needed to fine-tune the HARP model itself for a benchmark.
Benchmark Pb. size In training set? Nb. timeouts Nb. exceed res. best GF/s
2mm small yes 0 4 42.33
2mm medium no 2 7 96.44
2mm large no 10 N/A N/A
gemm small yes 0 0 27.38
gemm medium yes 0 0 125.59
gemm large no 3 5 10.65
gramsch. small no 0 0 1.16
gramsch. medium no 10 N/A N/A
gramsch. large no 10 N/A N/A
Table 2. Performance of HARP on double-precision floating-point data, for 2mm, Gemm and Gramschmidt varying the problem size
@@ 2.3. Limitations of General-Purpose DSE
AutoDSE (Sohrabizadeh et al., 2022b) utilizes the HLS compilers as a black box, in order to select the configurations that minimize the objective function. The tools are agnostic of the input program shape and if it detects that Merlin did not apply the pragmas as expected it allows the DSE to prune the design after Merlin has generated the HLS-C code. These frameworks use incremental DSEs, i.e., having no information on the characteristics of the program, they explore the space by increasing the parallelism in order to respond to a problem, e.g., a bottleneck for AutoDSE.
@@
Table 3. Investigation of design space and exploration coverage for synthesized, pruned, and timeout designs by the AutoDSE framework for the 2mm, Gemm, and Gramschmidt kernels.
Lack of full understanding of pragma impact
@@
Botlleneck analysis will make it possible to select which part of the code to optimize as a priority (Sohrabizadeh et al., 2022b). However, this priority does not take into account parallelism (i.e., hardware resources) that shall be deployed for other parts of the code. This creates code with regions that are extremely/fully parallelized, and others without any parallelism.
For 2mm, the fastest design found by AutoDSE mainly optimizes one loop body. When AutoDSE tries to optimize the second loop body, it favors the unroll factors to the power of two for the innermost loop and goes directly to the outermost loop. This does not enhance performance or generate configurations pruned by AutoDSE as the pragmas are not applied by AMD Merlin. The fact that it does not try the other unrolls factors for the innermost loop before optimizing the other pragmas leads a loss of performance. For Gemm and Gramschmidt, the DSE finds designs with a good QoR. However the DSE wastes much time exploring too large unroll factors, which generates ponderously long synthesis times without giving any result as the HLS timeout is reached. The time spent increasing the unroll factor for certain pragmas without result does not allow the unroll factor of other pragmas to be increased, which results in missed performance.
Limitations of filtering-based approaches
Approaches that rely on actual HLS runs for exploration have the benefit of being independent of the toolchain being used, as actual synthesis is used to obtain the QoR for designs. Accelerating the search by using a predictor, typically to filter out designs with low predicted QoR can significantly reduce search time and offer solid performance (Sohrabizadeh et al., 2023a). However, training requires a significant database of actual designs synthesized by HLS (e.g. 80,000 for HARP (Bai et al., 2024)), itself a particularly time-consuming process. Extending to other data types, other problem sizes, or even other benchmarks may require costly additional retraining or fine-tuning to achieve good performance. In addition, these models are “black-boxes” and do not provide meaningful insight of the predictions they produce. In contrast, our proposed NLP-DSE approach offers interpretable estimations, enabling developers to better understand the performance achieved, similar to traditional analytical models. It also guarantees that any pruned design is slower than the best design found so far. By leveraging the lower-bound based cost models, our approach estimates latency for specific configurations faster than neural network based approaches for affine programs, allowing for the exploration of many more designs within the same time frame. Moreover, given the explicit latency and resource functions we developed, we can employ existing non-linear programming (NLP) to directly identify the most promising candidate designs for HLS validation, significantly reduce the number of HLS runs.
3. Overview of NLP-DSE